OpenStudy (anonymous):

How long would it take a penny to hit the ground when droped from a roof 192 ft high

OpenStudy (alfie):

yf = yi +v0t+1/2at^2 yf = 0 yi = 192ft v0t = 0 1/2at^2 = -1/2gt^2 Where g = -9.8m/s^2

OpenStudy (alfie):

or well, since you are dealing with feet you should plug in the correct "g" value ^_^

OpenStudy (anonymous):

Huh

OpenStudy (anonymous):

Use -16ft/s^2 for this equation converting to meters is a waste of time if this isn't physics. he's saying since you a dropping it from rest it does not have an initial velocity. Use -g/2t^2+v0t+s0=s(t) as he was saying your initial height is represented by 192 ft in other words s0=192, your initial velocity, v0, is 0 so no need to plug anything in there, and finally your g is 16 so plug in 32 finally you end up with -32/2t^2+0+192=s(t) or -16t^2+0+192=s(t) we want to know the time when s=0 so since this equation is the distance equation solve set it equal to zero and solve for t that should be your answer.